# London WC2R 2LS, UK

### v1

@@ 2.2 Geometric $c$-theorem and monotonicity of renormalization group flow
where $\lambda=\lambda(G,B,A)$ is the eigenvalue of the ground state. A Schrödinger-like operator has a unique normalized positive-definite eigenfunction associated with the ground state which can be identified with $e^{-\Phi}=e^{-\Phi}(G,B,A)$. This ground state is the unique extremum of $S(G,B,A,\Phi)$ under the constraint $C(G,\Phi)=0$, and in particular it is the absolute maximum as $S(G,B,A,\Phi)|_{\Phi=\Phi(G,B,A),C=0}=-\lambda(G,B,A)$.
Thus the constrained maximum of $S(G,B,\Phi)$ as a function of $\Phi$ for fixed $G$, $B$ and $A$ always exists and is unique. Because of this uniqueness, and general properties of elliptic partial differential equations, the maximum varies smoothly as a function of $G$, $B$ and $A$. Therefore, $\overline{S}(G,B,A)\equiv S(G,B,A,\Phi(G,B,A))$, where $\Phi$ is regarded as a function $\Phi(G,B,A)$, evaluated at the constrained maximum of $S$, is a smooth function of $G$, $B$ and $A$.
@@
$\frac{\delta\overline{S}}{\delta A_{Iab}}$ $=e^{-2\widehat{\Phi}}\beta_{A}^{Iab}=-\frac{\alpha^{\prime}}{2}\widehat{\nabla}_{J}\left(e^{-2\widehat{\Phi}}F^{JIab}\right)~{},$ (24)
where $\widehat{\Phi}=\Phi-\log\sqrt[{}^{4}]{G}$, $\beta_{G}$, $\beta_{B}$ and $\beta_{A}$ are the heterotic sigma model beta functions and ^$\widehat{\nabla}_{J}Y^{Ia}={\mathcal{D}}_{J}Y^{Ia}+\widehat{\Gamma}^{I}_{JK}Y^{Ka}$. Therefore from (21), one concludes that the first variation of the functional $\overline{S}$ along $G$, $B$ and $A$ can be written as
@@ 2.3 Conformal invariance from scale invariance
00 0 $=\int_{M^{D}}{\mathrm{d}}^{D}Xe^{-2\widehat{\Phi}}\Big{(}\beta_{G}^{IJ}B_{G\,IJ}+\beta_{B}^{IJ}B_{B\,IJ}+\beta_{A}^{Iab}B_{A\,Iab}\Big{)}$ (47)
$=\int_{M^{D}}{\mathrm{d}}^{D}Xe^{-2\widehat{\Phi}}\left(\beta_{G}^{IJ}\beta_{G\,IJ}+\beta_{B}^{IJ}\beta_{B\,IJ}+\beta_{A}^{Iab}\beta_{A\,Iab}\right)\,,$ (48)
as the $\overline{S}$ is invariant under diffeomorphisms, $B$-field gauge transformations and the gauge transformations of $A$, i.e.
@@
irrespective on whether the beta functions vanish. Therefore from (47), we conclude that $\beta_{G}=\beta_{B}=\beta_{A}=0$ and all scale invariant heterotic sigma models are conformally invariant with target spaces compact manifolds $M^{D}$. The proof above applies to up to two loops in sigma model perturbation theory. Using a similar argument as that presented in PapWitten for the the common sector of string theory, it can be extended to all loops in perturbation theory.
@@ 3 Geometry, scale and conformal Invariance
To give some examples of sigma models that exhibit scale and conformal invariance, we shall exploit some special geometries which naturally arise in the context of heterotic geometry. Typically, these involve an appropriate restriction on the holonomy of the connection $\widehat{\nabla}$ on $M^{D}$. It turns out that there are several such geometries that can be found. In all cases, the associated heterotic sigma models admit additional symmetries generated by the $\widehat{\nabla}$-covariantly constant forms of these geometries. These include additional worldsheet supersymmetries as well as additional W-type of symmetries. For the class of such geometries which will be explored below, the holonomy of $\widehat{\nabla}$ on $M^{D}$ is restricted to $U(n)$ $(D=2n)$, $SU(n)$ $(D=2n)$, $Sp(k)$ $(D=4k)$, $G_{2}$ $(D=7)$ and $\mathrm{Spin}(7)$ $(D=8)$, see also (1) in the introduction.
3.1 KT and CYT geometries
The Kähler with torsion (KT) and the Calabi-Yau with torsion (CYT) geometries [^5] KT, CYT and HKT (see section ) geometries are also referred to in the literature, especially for bi-KT, bi-CYT and bi-HKT, as generalised Kähler, generalised Calabi-Yau and generalised hyper-Kähler, respectively Hitchin ; Gualtieri – bi-KT, bi-CYT and bi-HKT geometries contain two copies of the associated structure one with respect to the connection $\widehat{\nabla}$ and the other with respect to the connection $\breve{\nabla}$. Here, we use the terminology of HoweGP to describe these geometries. The strong version of these geometries is also referred to as pluriclosed. have recently been reviewed PapWitten in a related context focusing on the strong case that the 3-form is closed, ${\mathrm{d}}H=0$. Because of this, here we shall only summarise without proof some of their key properties. However, we shall emphasise the differences that arise in the properties of these geometries whenever $H$ is not closed, ${\mathrm{d}}H\not=0$, as this is the case that applies to the heterotic sigma models, see (12). To begin suppose that $M^{D}$, $D=2n$, is a KT (Kähler with torsion) manifold, i.e. a Hermitian manifold equipped with the unique connection $\widehat{\nabla}$ with skew-symmetric torsion [^6] The torsion $H=-\iota_{I}{\mathrm{d}}I=-{\mathrm{d}}_{I}I$, where ${\mathrm{d}}_{I}=\iota_{I}{\mathrm{d}}-{\mathrm{d}}\iota_{I}={\mathrm{i}}(\partial-\overline{\partial})$, $\partial$ is the holomorphic exterior derivative and the inner derivation $\iota_{I}L$ of a $k$-form $L$ with respect to $I$ is $\iota_{I}L=\frac{1}{(k-1)!}I^{K}{}_{P_{1}}L_{KP_{2}\dots P_{k}}{\mathrm{d}}x^{P_{1}}\wedge\dots\wedge{\mathrm{d}}x^{P_{k}}$. $H$ such that both the metric $G$ and complex structure $I$ are covariantly constant, ^$\widehat{\nabla}G=\widehat{\nabla}I=0$. As consequence of ^$\widehat{\nabla}G=\widehat{\nabla}I=0$ is that the holonomy of $\widehat{\nabla}$ reduces from $SO(2n)$ to $U(n)$. The 3-form $H$ is of type [^7] On complex manifolds, a $k$-form can be decomposed in components characterised by a bi-degree $(r,s)$, with $r+s=k$, where $r$ denotes the number of holomorphic and $s$ denotes the number of anti-holomorphic directions of the form. $(2,1)\oplus(1,2)$ with respect to the complex structure $I$, i.e. its (3,0) and (0,3) components vanish. This is a consequence of ^$\widehat{\nabla}I=0$ together with the integrability of the complex structure $I$.
As $\widehat{\nabla}$ is a metric connection and ^$\widehat{\nabla}I=0$, the curvature tensor $\widehat{R}$ of the connection $\widehat{\nabla}$ is skew-symmetric in the last two indices, $\widehat{R}_{PQST}=-\widehat{R}_{PQTS}$, and satisfies the (1,1) condition
@@ 3.1 KT and CYT geometries
This condition is a consequence the reduction of the holonomy of $\widehat{\nabla}$ from $SO(2n)$ to $U(n)$ and arises as the integrability condition of ^$\widehat{\nabla}I=0$. The Lee form of a KT manifold is defined as
@@
where the second equality follows as a consequence of ^$\widehat{\nabla}I=0$.
@@
$\widehat{R}_{IJ}=\widehat{\nabla}_{I}\widehat{\theta}_{J}+\frac{1}{4}dH_{ILPQ}I^{L}{}_{J}I^{PQ}\,.$ (55)
after using both (51) and (53) as well as ^$\widehat{\nabla}I=0$ and the definition of the Lee form (52).
Further progress depends on being able to reexpress the $dH$ term in the equation above in terms of quadratic terms in $\widetilde{R}$ and $F$ that appear in the beta function for the metric in (21). One way to achieve this is to assume that $M^{D}$ admits a connection $\widetilde{\nabla}$ such that
@@
and similarly for the connection [^8] These conditions (57) on a connection are a special case of Hermitian-Einstein or Hemitian-Yang-Mills equations and reduce on a 4-dimensional complex manifold to the usual anti-self-duality condition. Provided certain stability conditions are met, the Hermitian-Einstein equations have always a solution on any compact Kähler manifold as a consequence of Donalson-Uhlenbeck-Yau theorem Donaldson ; UhlenbeckYau and some classes of Hermitian manifolds LiYau , for a recent survey see Li . of vector bundle $E$
@@
Using conditions (56) and (57), the equation (55) can be re-arranged as
@@
The symmetric and antisymmetric parts of this equation are actually equivalent to the conditions (44) for global scale-invariance, with $V$ and $\Lambda$ being related to $\widehat{\theta}$. A detailed investigation of the conditions will be described in section 4.1.
To the order in perturbation theory we consider, one can identify $\widetilde{R}$ with $\breve{R}$ – this is the curvature of the connection $\breve{\nabla}$ that has torsion $-H$. To justify this choice, the Bianchi identity that relates the curvatures of $\widehat{\nabla}$ and $\breve{\nabla}$ is
@@
and contract this with $I^{K}{}_{I}I^{PQ}$. After expressing ${\mathcal{D}}$ in terms of the connection with torsion, using ^$\widehat{\nabla}I=0$ and the conditions (57), one finds that
^$\widehat{\nabla}^{J}F_{JI}{}^{a}{}_{b}+\widehat{\theta}^{J}F_{JI}{}^{a}{}_{b}=0~{}.$ (62)
@@ 3.2.1 HKT geometry
The Hyper-Kähler Geometry with Torsion (HKT) has also been reviewed in a similar context in PapWitten and so we shall only briefly explain the key points that will be useful later. An HKT manifold is a manifold of dimension $D=2n=4k$ endowed with three integrable complex structures $I_{r}$, $r=1,2,3$ that satisfy the algebra of imaginary unit quaternions: $I_{1}^{2}=I_{2}^{2}=-1$, $I_{1}I_{2}+I_{2}I_{1}=0$, and $I_{3}=I_{1}I_{2}$, or equivalently $I_{r}I_{s}=-\delta_{rs}1+\epsilon_{rs}{}^{t}I_{t}$. In addition, it is equipped with a metric $G$ and torsion $H$, which is a 3-form, such that the metric $G$ is Hermitian with respect to each of the complex structures $I_{r}$, $r=1,2,3$, and $I_{r}$ are covariantly constant, ^$\widehat{\nabla}I_{r}=0$, with respect to the connection $\widehat{\nabla}$. The geometry of the twistor space of an HKT manifolds and other properties have been investigated sometime ago in HoweGP . More recently, it has been shown that the moduli spaces of instantons on HKT manifolds admit an HKT structure RMMV following earlier work in LT ; NH . For a detailed description of the geometry of the moduli space of instantons on the HKT manifold $S^{3}\times S^{1}$ as well as its applications in AdS/CFT see Witten .
@@
This follows from the integrability of the complex structures and ^$\widehat{\nabla}I_{r}=0$. Also, it is valid irrespectively on whether $H$ is closed or not.
As $\widehat{\nabla}$ is a metric connection and ^$\widehat{\nabla}I_{r}=0$, the curvature $\widehat{R}$ of $\widehat{\nabla}$ is shew-symmetric in the last two indices and
@@
It turns out that this implies (53) for each of the complex structures $I_{r}$. The covariant constancy of $G$ and of the $I_{r}$ implies that the holonomy of $\widehat{\nabla}$ is contained in $Sp(k)\subset SU(n)\subset SO(D)$
@@ 3.2.2 $G_{2}$ and ${\mathrm{S}pin}(7)$ geometries with torsion
It is clear from the geometries investigated above that one way to construct solutions to the scale invariance conditions in (44) is to appropriately restrict the holonomy of $\widehat{\nabla}$. Apart from the two cases investigated so far, for which the holonomy of $\widehat{\nabla}$ is restricted to $SU(n)$ and $Sp(k)$, respectively, there are two more possibilities that can be explored. One is in dimension seven, $D=7$, for which the holonomy of $\widehat{\nabla}$ is restricted in $G_{2}\subset SO(7)$ and another in dimension eight, $D=8$, for which the holonomy of $\widehat{\nabla}$ is restricted in ${\mathrm{S}pin}(7)\subset SO(8)$.
To begin from the $G_{2}$ case, the $G_{2}$ geometry with torsion $H$, a 3-form, on $M^{7}$ with metric $G$ is characterised by the existence of a 3-form $\varphi$, the fundamental $G_{2}$ form, such that ^$\widehat{\nabla}\varphi=0$ tfsi1 ; tfsi2 , see also gmpw . In an adapted frame $({\mathbf{e}}^{A}={\mathbf{e}}^{i}_{I}dx^{I};A=1,\dots,7)$, where $x$ are some coordinates on the manifold, the metric and fundamental $G_{2}$ form $\varphi$ can be chosen as
@@
where ${\mathbf{e}}^{123}$ is a shorthand for ${\mathbf{e}}^{1}\wedge{\mathbf{e}}^{2}\wedge{\mathbf{e}}^{3}$ and similarly for the rest of the components of $\varphi$. The condition [^9] The same condition ^$\widehat{\nabla}\varphi=0$ also determines $H$ in terms of $G$ and the fundamental form $\varphi$ and its first derivatives as for KT geometries. ^$\widehat{\nabla}\varphi=0$ implies that the curvature $\widehat{R}$ satisfies that
@@
and $\widehat{R}_{PQKL}=-\widehat{R}_{PQLK}$ as $\widehat{\nabla}$ is metric, where ${}^{*}\varphi$ is the dual form of $\varphi$ on $M^{7}$. To see this, the integrability condition of ^$\widehat{\nabla}\varphi=0$ implies that
@@
on the curvatures $\widetilde{R}$ and $F$ of the connection $\widetilde{\nabla}$ and ${\mathcal{D}}$ – we shall refer to these equations as $G_{2}$ instanton-like conditions. Contracting the Bianchi identity (54) with ${}^{*}\varphi^{LPQ}{}_{T}$ and using the above equations, one confirms that the $G_{2}$ geometry with torsion satisfies the same scale invariance condition (58) as the KT geometry with Lee form $\widehat{\theta}$
@@
where in the last step ^$\widehat{\nabla}\varphi=0$ has been used as well as (71). Again at the order in perturbation theory we are considering, one can set $\widetilde{R}=\breve{R}$ as in the KT case.
@@
${}^{*}\varphi^{LIJK}{}^{*}\varphi_{LPQR}=6\,\delta^{I}_{[P}\delta^{J}_{Q}\delta^{K}_{R]}-9\,\delta^{[I}_{[P}\,{}^{*}\varphi^{JK]}{}_{QR]}~{},$ (74)
one can verify that $F$ satisfies (62).
Similarly to the $G_{2}$ geometry with torsion, the ${\mathrm{S}pin}(7)$ geometry with torsion $H$, a 3-form, on $M^{8}$ with metric $G$ is characterised by the existence of a self-dual 4-form $\phi$, the fundamental ${\mathrm{S}pin}(7)$ form, such that ^$\widehat{\nabla}\phi=0$ si2 . In an adapted frame $({\mathbf{e}}^{A}={\mathbf{e}}^{A}_{I}dx^{I};A=1,\dots,8)$ on $M^{8}$, the metric and fundamental form on $M^{8}$ can be written as
@@
The condition [^10] The same condition ^$\widehat{\nabla}\phi=0$ also determines $H$ in terms of $G$ and the fundamental form $\phi$ and its first derivatives as for KT geometries. on the curvature $\widehat{R}$ of $\widehat{\nabla}$ that arises as an integrability of ^$\widehat{\nabla}\phi=0$ can be expressed as
@@
where the curvature is skew-symmetric in the last two indices as $\widehat{\nabla}$ is metric. This can be shown following similar steps to those used to prove (68) and the identity
@@
on the curvature of the $\widetilde{\nabla}$ connection and that of the gauge sector. Under these conditions one can again establish that this geometry satisfies the same scale invariance condition as that of the KT geometry (58) but now with Lee form
@@
To establish the second part of the identity one uses ^$\widehat{\nabla}\phi=0$ and (78). In the order of perturbation theory we are considering, one can again set $\widetilde{R}=\breve{R}$.

### v2

@@ 2.2 Geometric $c$-theorem and monotonicity of renormalization group flow
where $\lambda=\lambda(G,B,A)$ is the eigenvalue of the ground state. The above Schrödinger-like operator has a unique normalized positive-definite eigenfunction associated with the ground state which can be identified with $e^{-\Phi}=e^{-\Phi}(G,B,A)$. This ground state is the unique extremum of $S(G,B,A,\Phi)$ under the constraint $C(G,\Phi)=0$, and in particular it is the absolute maximum as $S(G,B,A,\Phi)|_{\Phi=\Phi(G,B,A),C=0}=-\lambda(G,B,A)$.
Thus the constrained maximum of $S(G,B,\Phi)$ as a function of $\Phi$ for fixed $G$, $B$ and $A$ always exists and it is unique. Because of this uniqueness, and general properties of elliptic partial differential equations, the maximum varies smoothly as a function of $G$, $B$ and $A$. Therefore, $\overline{S}(G,B,A)\equiv S(G,B,A,\Phi(G,B,A))$ evaluated at the constrained maximum of $S$, where $\Phi$ is regarded as a function $\Phi(G,B,A)$, is a smooth function of $G$, $B$ and $A$.
@@
$\frac{\delta\overline{S}}{\delta A_{Iab}}$ $=e^{-2\widehat{\Phi}}\beta_{A}^{Iab}=\frac{\alpha^{\prime}}{2}\widehat{\nabla}_{J}\left(e^{-2\widehat{\Phi}}F^{JIab}\right)~,$ (24)
where $\widehat{\Phi}=\Phi-\log\sqrt[{}^{4}]{G}$, $\beta_{G}$, $\beta_{B}$ and $\beta_{A}$ are the heterotic sigma model beta functions and $\widehat{\nabla}_{J}Y^{Ia}={\mathcal{D}}_{J}Y^{Ia}+\widehat{\Gamma}^{I}_{JK}Y^{Ka}$. Therefore from (21), one concludes that the first variation of the functional $\overline{S}$ along $G$, $B$ and $A$ can be written as
@@ 2.3 Conformal invariance from scale invariance
$\int_{M^{D}}{\mathrm{d}}^{D}Xe^{-2\widehat{\Phi}}\Big{(}\beta_{G}^{IJ}B_{G\,IJ}+\beta_{B}^{IJ}B_{B\,IJ}+\beta_{A}^{Iab}B_{A\,Iab}\Big{)}=0\,.$ (47)
But $\overline{S}$ is invariant under diffeomorphisms, $B$-field gauge transformations and the gauge transformations of $A$ and so
@@
irrespective on whether the beta functions vanish. Using this and (47), we conclude that
$\int_{M^{D}}{\mathrm{d}}^{D}Xe^{-2\widehat{\Phi}}\left(\beta_{G}^{IJ}\beta_{G\,IJ}+\beta_{B}^{IJ}\beta_{B\,IJ}+\beta_{A}^{Iab}\beta_{A\,Iab}\right)=0\,$, (50)
and so $\beta_{G}=\beta_{B}=\beta_{A}=0$. Thus all scale invariant heterotic sigma models are conformally invariant with target spaces compact manifolds $M^{D}$. The proof above applies up to two loops in sigma model perturbation theory. Using a similar argument as that presented in PapWitten for the the common sector of string theory, it can be extended to all loops in perturbation theory.
@@ 3 Geometry, scale and conformal Invariance
To give some examples of sigma models that exhibit scale and conformal invariance, we shall exploit some special geometries which naturally arise in the context of heterotic geometry. Typically, these involve an appropriate restriction on the holonomy of the connection $\widehat{\nabla}$ on $M^{D}$. It turns out that there are several such geometries that can be found. In all cases, the associated heterotic sigma models admit additional symmetries generated by the $^\widehat{\nabla}$-covariantly constant forms of these geometries. These include additional worldsheet supersymmetries as well as additional W-type of symmetries. For the class of such geometries that will be explored below, the holonomy of $\widehat{\nabla}$ on $M^{D}$ is restricted to $U(n)$ $(D=2n)$, $SU(n)$ $(D=2n)$, $Sp(k)$ $(D=4k)$, $G_{2}$ $(D=7)$ and $\mathrm{Spin}(7)$ $(D=8)$, see also (1) in the introduction.
3.1 Scale invariance, KT and CYT geometries
[…]
As $\widehat{\nabla}$ is a metric connection and $\widehat{\nabla}I=0$, the curvature tensor $\widehat{R}$ of the connection $\widehat{\nabla}$ is skew-symmetric in the last two indices, $\widehat{R}_{PQST}=-\widehat{R}_{PQTS}$, and satisfies the (1,1) condition
@@ 3.1 Scale invariance, KT and CYT geometries
This condition is a consequence the reduction of the holonomy of $\widehat{\nabla}$ from $SO(2n)$ to $U(n)$ and arises as the integrability condition of $\widehat{\nabla}I=0$. The Lee form of a KT manifold is defined as
@@
where the second equality follows because of $\widehat{\nabla}I=0$.
@@
$\widehat{R}_{IJ}=\widehat{\nabla}_{I}\widehat{\theta}_{J}+\frac{1}{4}dH_{ILPQ}I^{L}{}_{J}I^{PQ}\,$, (55)
where we have used both (51) and (53) as well as $\widehat{\nabla}I=0$ and the definition of the Lee form (52).
Further progress depends on being able to reexpress (55) as the beta function for the metric in (21) and in particular the last term of the former equation as the quadratic expression in the curvatures $\widetilde{R}$ and $F$ that appears in the latter equation. For this, one can use the expression for $dH$ in (12) and assume that $M^{D}$ admits a connection $\widetilde{\nabla}$ such that
@@
and similarly for the connection [^9] These conditions (57) on the connection $A$ are a special case of Hermitian-Einstein or Hermitian-Yang-Mills equations and reduce on a 4-dimensional complex manifold to the usual anti-self-duality condition. Provided certain stability conditions are met, the Hermitian-Einstein equations always admit a solution on any compact Kähler manifold as a consequence of Donalson-Uhlenbeck-Yau theorem Donaldson ; UhlenbeckYau and some classes of Hermitian manifolds LiYau , for a recent survey see Li . of vector bundle $E$
@@
Using conditions (56), (57) and (12), the equation (55) can be re-arranged as
@@
The symmetric and antisymmetric parts of this equation are actually equivalent to the conditions (44) for global scale-invariance, with $V$ and $\Lambda$ being related to $\widehat{\theta}$. A detailed investigation of these conditions will be given in section 4.1.
[…]
@@
and contract this with $I^{K}{}_{I}I^{PQ}$. After expressing $\mathcal{D}$} in terms of the connection with torsion, using $\widehat{\nabla}I=0$ and the conditions (57), one finds that
$\widehat{\nabla}^{J}F_{JI}{}^{a}{}_{b}+\widehat{\theta}^{J}F_{JI}{}^{a}{}_{b}=0~.$ (62)
@@ 3.2.1 HKT geometry
The Hyper-Kähler Geometry with Torsion (HKT) has also been reviewed in a similar context in PapWitten and so we shall only briefly explain the key points that will be useful later. An HKT manifold is a manifold of dimension $D=2n=4k$ endowed with three integrable complex structures $I_{r}$, $r=1,2,3$ that satisfy the algebra of imaginary unit quaternions: $I_{1}^{2}=I_{2}^{2}=-1$, $I_{1}I_{2}+I_{2}I_{1}=0$, and $I_{3}=I_{1}I_{2}$, or equivalently $I_{r}I_{s}=-\delta_{rs}{\mathbf{1}}+\epsilon_{rs}{}^{t}I_{t}$. In addition, it is equipped with a metric $G$ and torsion $H$, which is a 3-form, such that the metric $G$ is Hermitian with respect to each of the complex structures $I_{r}$, $r=1,2,3$, and $I_{r}$ are covariantly constant, $\widehat{\nabla}I_{r}=0$, with respect to the connection $\widehat{\nabla}$. The geometry of the twistor space of an HKT manifolds and other properties have been investigated sometime ago in HoweGP . More recently, it has been shown that the moduli spaces of instantons on HKT manifolds admit an HKT structure RMMV following earlier work in LT ; NH . For a detailed description of the geometry of the moduli space of instantons on the HKT manifold $S^{3}\times S^{1}$ as well as its applications in AdS/CFT see Witten .
@@
This follows from the integrability of the complex structures and $\widehat{\nabla}I_{r}=0$. Also, it is valid irrespectively on whether $H$ is closed or not.
As $\widehat{\nabla}$ is a metric connection and $\widehat{\nabla}I_{r}=0$, the curvature $\widehat{R}$ of $\widehat{\nabla}$ is shew-symmetric in the last two indices and
@@
It turns out that this implies (53) for each of the complex structures $I_{r}$. The covariant constancy of $G$ and that of $I_{r}$ implies that the holonomy of $\widehat{\nabla}$ is contained in $Sp(k)\subset SU(2k)\subset SO(D)$
@@ 3.2.2 $G_{2}$ and ${\mathrm{S}pin}(7)$ geometries with torsion
It is clear from the geometries investigated above that one way to construct solutions to the scale invariance conditions in (44) is to appropriately restrict the holonomy of $\widehat{\nabla}$. Apart from the two cases investigated so far, for which the holonomy of $\widehat{\nabla}$ is restricted to $SU(n)$ and $Sp(k)$, respectively, there are two more possibilities that can be explored. One is in dimension seven, $D=7$, for which the holonomy of $\widehat{\nabla}$ is restricted to be in $G_{2}\subset SO(7)$ and another in dimension eight, $D=8$, for which the holonomy of $\widehat{\nabla}$ is restricted to be in ${\mathrm{S}pin}(7)\subset SO(8)$.
To begin from the $G_{2}$ case, the $G_{2}$ geometry with torsion $H$, a 3-form, on $M^{7}$ with metric $G$ is characterised by the existence of a 3-form $\varphi$, the fundamental $G_{2}$ form, such that $\widehat{\nabla}\varphi=0$ tfsi1 ; tfsi2 , see also gmpw . In an adapted frame $({\mathbf{e}}^{A}={\mathbf{e}}^{i}_{I}dx^{I};A=1,\dots,7)$, where $x$ are some coordinates on the manifold, the metric and fundamental $G_{2}$ form $\varphi$ can be chosen as
@@
where ${\mathbf{e}}^{123}$ is a shorthand for ${\mathbf{e}}^{1}\wedge{\mathbf{e}}^{2}\wedge{\mathbf{e}}^{3}$ and similarly for the rest of the components of $\varphi$. The condition [^10] As for KT geometries, the condition $\widehat{\nabla}\varphi=0$ determines $H$ in terms of $G$, the fundamental form $\varphi$ and its first derivatives. $\widehat{\nabla}\varphi=0$ implies that the curvature $\widehat{R}$ satisfies that
@@
and $\widehat{R}_{PQKL}=-\widehat{R}_{PQLK}$ as $\widehat{\nabla}$ is a metric connection, where $^{*}\varphi$ is the dual form of $\varphi$ on $M^{7}$. To see this, the integrability condition of $\widehat{\nabla}\varphi=0$ implies that
@@
on the curvatures $\widetilde{R}$ and $F$ of the connection $\widetilde{\nabla}$ and $\mathcal{D}$} – we shall refer to these equations as $G_{2}$ instanton-like conditions. Contracting the Bianchi identity (54) with $^{*}\varphi^{LPQ}{}_{T}$ and using the above equations and (12), one confirms that the $G_{2}$ geometry with torsion satisfies the same scale invariance condition (58) as the CYT geometry with Lee form $\widehat{\theta}$
@@
where in the last step $\widehat{\nabla}\varphi=0$ has been used as well as (71). Again at the order in perturbation theory we are considering, one can set $\widetilde{R}=\breve{R}$ as in the CYT case.
@@
$^{*}\varphi^{LIJK}{}^{*}\varphi_{LPQR}=6\,\delta^{I}_{[P}\delta^{J}_{Q}\delta^{K}_{R]}-\varphi^{IJK}\varphi_{PQR}-9\,\delta^{[I}_{[P}\,{}^{*}\varphi^{JK]}{}_{QR]}~$, (74)
one can verify that $F$ satisfies (62) with Lee form $\widehat{\theta}$ given in (73).
Similarly to the $G_{2}$ geometry with torsion, the ${\mathrm{S}pin}(7)$ geometry with torsion $H$, a 3-form, on $M^{8}$ with metric $G$ is characterised by the existence of a self-dual 4-form $\phi$, the fundamental ${\mathrm{S}pin}(7)$ form, such that $\widehat{\nabla}\phi=0$ si2 . In an adapted frame $({\mathbf{e}}^{A}={\mathbf{e}}^{A}_{I}dx^{I};A=1,\dots,8)$ on $M^{8}$, the metric and fundamental form on $M^{8}$ can be written as
@@
The condition [^11] The same condition $\widehat{\nabla}\phi=0$ also determines $H$ in terms of $G$ and the fundamental form $\phi$ and its first derivatives as for KT geometries. on the curvature $\widehat{R}$ of $\widehat{\nabla}$ that arises as an integrability of $\widehat{\nabla}\phi=0$ can be expressed as
@@
where the curvature is skew-symmetric in the last two indices as $\widehat{\nabla}$ is a metric connection. This can be shown following similar steps to those used to prove (68) and by utilising the identity
@@
on the curvature of the $\widetilde{\nabla}$ connection and that of the gauge sector. Under these conditions one the curvatures and (12), one can again establish that this geometry satisfies the same scale invariance condition as that of the CYT geometry (58) but now with Lee form
@@
To establish the second part of the identity above, one uses $\widehat{\nabla}\phi=0$ and (78). In the order of perturbation theory we are considering, one can again set $\widetilde{R}=\breve{R}$.
@@
$\mathrm{Geometry}$ $\mathrm{Holonomy}$ $\mathrm{Scale~Invariant}$ $\mathrm{Conformal~Invariant}$
$\mathrm{KT}$ $U(n)$, $D=2n$ $\mathrm{No}$ $\mathrm{No}$
$\mathrm{CYT}$ $SU(n)$, $D=2n$ $\mathrm{Yes}$ $\mathrm{Yes}$
$\mathrm{HKT}$ $Sp(k)$, $D=4k$ $\mathrm{Yes}$ $\mathrm{Yes}$
$G_{2}$ $\mathrm{with~torsion}$ $G_{2}$, $D=7$ $\mathrm{Yes}$ $\mathrm{Yes}$
$\mathrm{Spin}(7)$ $\mathrm{with~torsion}$ $\mathrm{Spin}(7)$, $D=8$ $\mathrm{Yes}$ $\mathrm{Yes}$
Table 1: Heterotic sigma models with target space a manifold with KT geometry are neither scale nor conformally invariant. Those with target space a CYT, HKT, $G_{2}$ or $\mathrm{Spin}(7)$ manifold with torsion are scale invariant up to and including two loops in sigma model perturbation theory provided that the gauge sector connection is chosen to be a Hermitian-Einstein, or equivalently $SU(n)$ instanton, $Sp(k)$ instanton, $G_{2}$ instanton or $\mathrm{Spin}(7)$ instanton connection, respectively, and the anomalous Bianchi identity holds. These scale invariant sigma models are also conformally invariant provided that the target space $M^{D}$ is compact. There are explicit geometries that illustrate these results in all cases.